# Role of Noise in the Fairen-Velarde model of bacterial respiration

### v1

@@ 5 Summary and concluding remarks
The Fairen-Velarde model is a prototypical model of bacterial respiration. The two-dimensional deterministic nonlinear differential equation for the oxygen and the nutrient contents can grasp the essential features of bacterial respirations. Depending on the various domains of the parameters, it can successfully provide the active phase (limit cycle) and the inactive phase (fixed point) of bacterial respiration. The existence of two time scales before reaching the fixed point has been shown in our previous work published in Int. J. Mod. Phys C (2024) 2450097.
In this article, we mainly focused on the role of noise in the steady behaviour of bacterial respiration. The stochasticity (introduced by noise) in the modified Fairen-Velrade model gives rise to various interesting phenomena. We have observed that the noise accelerates the motion towards the fixed point. We have shown the reduction of the gross time scale due to the presence of noise. This has far-reaching consequences in bacteriology. Just by introducing the noise (which may be thermal, impurity, or mechanical noise), one can bring the system of bacteria into an inactive phase (fixed point) faster than in a noiseless case.
The fixed point remains no longer a fixed point in the presence of noise. The scattering of fixed-point data covers a fixed-point domain in the presence of noise. The area of such a domain has been found to increase as the noise intensity is increased. Asymptotically, this growth is linear with noise intensity. To reach the fixed point, it becomes probabilistic with a given time in the presence of noise. The probality of reaching the fixed point domain within a given interval of time shows a hyperbolic tangent behaviour with noise intensity.
@@ Acknowledgements:
Soumyadeep Kundu thankfully acknowledges the library facility of Presidency University, Kolkata.

### v3

@@ 5 Summary and concluding remarks
The Fairen-Velarde model is a prototypical model of bacterial respiration. The two-dimensional deterministic nonlinear differential equation for the oxygen and the nutrient contents can grasp the essential features of bacterial respirations. Depending on the various domains of the parameters, it can successfully provide the active phase (limit cycle) and the inactive phase (fixed point) of bacterial respiration. The existence of two time scales before reaching the fixed point has been shown in our previous work published in Int. J. Mod. Phys C, 35 (2024) 2450094.
In this article, we primarily focus on the role of noise in the steady-state behavior of bacterial respiration. Stochastic effects, introduced by noise in the modified Fairén-Velarde model, lead to several interesting phenomena. We observe that noise accelerates the system’s motion toward the fixed point, effectively reducing the gross time scale required to reach it. This finding has significant implications in bacteriology, as introducing noise—whether thermal, impurity-based, or mechanical—can expedite the transition of a bacterial system into the inactive phase (fixed point) compared to the noiseless case.
However, the presence of noise alters the nature of the fixed point. Instead of a single fixed point, the scattering of data creates a domain of fixed points. The size of this domain increases with the intensity of the noise, growing asymptotically in a linear manner. Additionally, in the presence of noise, reaching the domain of fixed points becomes probabilistic over a given time frame. The probability of reaching this domain within a specified time exhibits a hyperbolic tangent-like behavior as a function of noise intensity.
@@ Acknowledgements:
Soumyadeep Kundu thankfully acknowledges Md. Samsul Habib Shah, IIT Madras, for having valuable discussions about the principal component analysis method. We thank the anonymous reviewer for important suggestions.